# Network Oblivious Transfer via Noisy Channels: Limits and Capacities

### v2

@@ V-A2 OT over a point-to-point noisy channel
[37] A protocol $\Pi$ securely computes $\binom{2}{1}-\text{OT}^{k}$ perfectly if and only if for every pair of algorithms $\bar{A}=(\bar{A}_{1},\bar{A}_{2})$ that is admissible for protocol $\Pi$ and for all inputs $(X(m_{0},m_{1}),Z)$ and auxiliary input $\mathbf{C}$ (representing the total information transmitted over the noiseless public channel), $\bar{A}$ produces outputs $(U,V)$ such that the following conditions are satisfied:
@@
    $I(U^{\prime};Z^{\prime}|\mathbf{C},Z)=0,\,\,\,\text{and}\,\,\,I(U^{\prime};V|\mathbf{C},Z,Z^{\prime},M_{Z^{\prime}})=0.$
@@
Consider a noisy DMC between Alice and Bob. OT with perfect secrecy (unconditional security) is impossible over a DM-MAC if the players are unbounded cheaters. Also, OT is impossible even if only one of the players is an unbounded cheater.
@@ V-A3 OT over a DM-MAC
    $I(X_{i};V|\mathbf{C},Z_{i},Z^{\prime}_{i},M_{iZ^{\prime}_{i}})_{i\in\{1,2\}}$ $=0.$
@@ V-B OT over the DM-MAC
As previously discussed, from Alice’s perspective, SU-SBC resembles the uniform input to a binary MAC. On the other hand, from Bob’s perspective, it corresponds to the output of a distinguishable mixture of three channels: the complete erasure channel with probability $(1-p)^{2}$, a partial erasure channel with probability 2p(1p)2p(1-p) and a DM-MAC with probability $p^{2}$.
@@
V-C Secret Key Agreement over MAC by Public Discussion
Secret key agreement over a simple point-to-point channel is initially addressed in [42] and [7]. In [42], the author considers that Alice aims to share a secret key with Bob in the presence of a passive wiretapper. A noiseless one-way channel assists the main noisy channel with unlimited capacity. Also, it is assumed that Eve can receive all messages sent over the public channel without error, but the wiretapper can not change the messages without being detected.
@@ V-C Secret Key Agreement over MAC by Public Discussion
A generalization of the above setting can be found in [43, 44]. Also, the problem of secret key agreement over MAC is studied in [45], in which the authors assumed that each transmitter plays the role of wiretapper for the other transmitter. This is the MAC with confidential message [46]. Consider the following settings: Alice-1 and Alice-2 aim to share secret keys with Bob in the presence of a passive wiretapper. Alice-$i$, Bob, and Eve govern the input $X_{i}$, the output $Y$, and $E$, respectively. Furthermore, a public channel with unlimited capacity is available for one-way communication from the senders to Bob. Both resources (MAC and the public channel) are available for communication, but not at the same time. First, the senders are allowed to use the MAC $n$ times, and then they can use the public channel only once.
@@ VI-A Noiseless Binary Adder Channel (BAC)
Consider the following OT setting over the noiseless BAC [32]: two senders, Alice-1 and Alice-2, each hold two binary strings. Bob must select one string from each sender, with the unchosen strings remaining hidden. Communication occurs via the BAC, where the channel output is given by the sum of the inputs, $Y=X_{1}+X_{2}$. This channel deterministically reveals the inputs when they are equal (0 or $1$), but produces an erasure symbol ($e$) when they differ, making the output ambiguous. Erasures thus occur in two of the four possible input pairs.
@@ VII Conclusion
We investigated bounds on the oblivious transfer capacity of noisy MAC, focusing on secure multiparty computations involving two non-colluding senders and a single receiver. A protocol was proposed that is both correct and secure against honest-but-curious parties. Furthermore, we demonstrated that the protocol remains correct and secure when the receiver is dishonest, although the exact capacity in such scenarios remains undetermined. The OT capacity region for honest-but-curious players was established for a special reduced version with optimal rates, while for a malicious receiver, the study delineates a feasible rate region. The work delves into noisy MACs comprehensively, introducing reductions to symmetric basic correlations (SBCs) as defined in [6] and extending OT protocols to multi-sender configurations. Key contributions include precisely characterizing OT capacity for honest-but-curious settings, leveraging information-theoretic bounds, and significant progress in addressing adversarial behavior.
An intriguing future direction is the study of OT over MACs involving colluding parties. If the senders can collude with the receiver—a scenario closer to practical settings—the secrecy criteria become substantially stricter. Also, In [6], the authors offer a protocol secure and correct against a malicious sender when the correlation is slightly unfair, albeit no longer achieving positive rates. In the MAC setting, if we assume both senders are cheating, it can be conjectured that a similar protocol does not provide positive rates. But what about if just one of the senders is cheating? These challenging cases form the basis of our next research endeavors.
@@ Acknowledgment
The authors hereby acknowledge Andreas Winter, Rémi A. Chou, and Pin-Hsun Lin for valuable discussions during this research project.
@@ Appendix B Proof of Lemma 2
Let $Z:=\sum_{i=0}^{n-1}Z_{i}$. Then $Z$ is a sum of $n$ i.i.d. bounded random variables in $0,\log(|\mathcal{X}|)]$. Hoeffding’s inequality [41] states that for independent $Z_{i}\in[a,b]$:
@@ Appendix C Proof of Lemma 6
This lemma is an extended version of [6, Lemma 5]. It evaluates the likelihood of passing a ”typicality test” when input sequences are fed into a noisy MAC and output sequences are generated. The lemma bounds the probability that the channel’s output remains consistent with a typical input-output relationship, even under deviations in the input strings (e.g., due to cheating). In simpler terms, this lemma demonstrates that if someone attempts to manipulate the inputs or outputs, the probability of such manipulations going undetected is exponentially small under specific conditions.
@@
Using the Chernoff bound [6, Lemma 4] completes the proof. ∎
@@ Appendix D Proof of Theorem 9
$H(M_{\overline{Z^{\prime}}}|V,Z,Z^{\prime},M_{Z^{\prime}})=H(M_{\overline{Z^{\prime}}}).$ (41)
@@
This implies that given the information $V$, the choice $Z$, the effective choice $Z^{\prime}$ and $M_{Z^{\prime}}$, Bob has no additional information about $M_{\overline{Z^{\prime}}}$. Consider the mutual information between Alice’s bits and the information received by Bob $I(X;V)=I(M_{0},M_{1};V)$. Since Bob should learn $M_{Z^{\prime}}$ and nothing about $M_{\overline{Z^{\prime}}}$, we have:
$I(M_{Z^{\prime}},M_{\overline{Z^{\prime}}};V|Z,Z^{\prime},M_{Z^{\prime}})=$ $I(M_{Z^{\prime}};V|Z,Z^{\prime},M_{Z^{\prime}})+I(M_{\overline{Z^{\prime}}};V|Z,Z^{\prime},M_{Z^{\prime}}).$ (42)
@@
The first term should not be equal to zero because $V$ must also convey the value of $M_{Z^{\prime}}$ to Bob. The second term should be zero due to the perfect secrecy of OT. So consider the second term:
$I(M_{\overline{Z^{\prime}}};V|Z,Z^{\prime},M_{Z^{\prime}})$ $=I(M_{\overline{Z^{\prime}}};V|Z,Z^{\prime})$
$=H(M_{\overline{Z^{\prime}}}|Z,Z^{\prime})-H(M_{\overline{Z^{\prime}}}|V,Z,Z^{\prime})$
$=H(M_{\overline{Z^{\prime}}}|Z,Z^{\prime})-H(M_{\overline{Z^{\prime}}})$
$=-I(M_{\overline{Z^{\prime}}};Z,Z^{\prime}),$ (43)
@@
where the first equality is happened due the fact that $M_{\overline{Z^{\prime}}}$ and $M_{Z^{\prime}}$ are independent, and the third equality follows from (41).
@@
As $I(M_{\overline{Z^{\prime}}};Z,Z^{\prime})$ is a non-negative quantity, then we proved that $I(M_{\overline{Z^{\prime}}};Y|M_{Z^{\prime}},Z,Z^{\prime})$ is not a positive quantity. It is impossible unless $I(M_{\overline{Z^{\prime}}};Z,Z^{\prime})=0$, which implies that all inputs $(X,Z)$ are independent (We already clarified that this is often unattainable and in the most general case, we assume there is a known dependency between the inputs). Combining (42) and (D), we have:
$I(M_{Z^{\prime}},M_{\overline{Z^{\prime}}};V|Z,Z^{\prime},M_{Z^{\prime}})$ $=I(M_{Z^{\prime}};V|Z,Z^{\prime})$
$=H(M_{Z^{\prime}}|Z,Z^{\prime})-H(M_{Z^{\prime}}|V,Z,Z^{\prime})$
$=H(M_{Z^{\prime}}|Z,Z^{\prime}),$ (44)
where the second equality is since given $V$ and $Z^{\prime}$, the uncertainty about $M_{Z^{\prime}}$ is zero (correctness criterion). The term $H(M_{Z^{\prime}}|Z,Z^{\prime})$ is not equal to zero because we already assumed that $M_{Z^{\prime}}$ and $Z^{\prime}$ are independent. This is a contradiction. ∎
@@ Appendix E Proof of Theorem 10
The proof follows from an extended version presented in [37]. Consider the intuitive security criterion for Bob: if Alice-$i$ acts maliciously, she sends $X^{\prime}_{i}$ instead of $X_{i}$ (This means that Alice-$i$ can replace some bits with any deterministic or random bits in her flavor so that the Hamming distance between the original sequence and the changed sequence could be in the range $0\leq d_{H}\leq n$). In such a case, security constraint implies that the mutual information between Bob and Alice-$i$’s effective input given $X_{i}$ and $\mathbf{C}$ should be zero $(X^{\prime}_{i}\rightarrow(\mathbf{C},X_{i})\rightarrow Y)$. On the other hand, the mutual information between Bob’s final view $(V,Y)$ and Alice-$i$’s final view given $(\mathbf{C},X_{i},X^{\prime}_{i})$ should be zero $(U_{i}\rightarrow(\mathbf{C},X_{i},X^{\prime}_{i})\rightarrow(V,Y))$:
$I(Y;X^{\prime}_{i}|\mathbf{C},X_{i})=0,\qquad\text{and}\qquad I(V,Y;U_{i}|\mathbf{C},X_{i},X^{\prime}_{i})=0,$
where $Y=Z_{i}$ and $V=M^{\prime}_{iZ_{i}}$. Combining both, we have:
$I(Z_{i};X^{\prime}_{i}|$ $\mathbf{C},X_{i})+I(M^{\prime}_{iZ_{i}},Z_{i};U_{i}|\mathbf{C},X_{i},X^{\prime}_{i})$
$\,=I(Z_{i};X^{\prime}_{i}|\mathbf{C},X_{i})+I(Z_{i};U_{i}|\mathbf{C},X_{i},X^{\prime}_{i})+I(M^{\prime}_{iZ_{i}};U_{i}|\mathbf{C},Z_{i},X_{i},X^{\prime}_{i})$
$\stackrel{{\scriptstyle(a)}}{{=}}I(Z_{i};X^{\prime}_{i}|\mathbf{C},X_{i})+I(Z_{i};U_{i}|\mathbf{C},X_{i},X^{\prime}_{i})$
$\,=I(Z_{i};X^{\prime}_{i},U_{i}|\mathbf{C},X_{i})$
$\,=I(Z_{i};U_{i}|\mathbf{C},X_{i})+I(Z_{i};X^{\prime}_{i}|\mathbf{C},X_{i},U_{i})$
$\stackrel{{\scriptstyle(b)}}{{=}}I(Z_{i};U_{i}|\mathbf{C},X_{i})$
@@
where $(a)$ follows from the fact that $I(M^{\prime}_{iZ_{i}};U_{i}|\mathbf{C},Z_{i},X_{i},X^{\prime}_{i})=0$, and $(b)$ is due to the following argument: choose $M^{\prime}_{iZ_{i}}=(M^{\prime}_{i,0},M^{\prime}_{i,1})$ and for the values $l\in\{0,1\}$, $M^{\prime}_{i,l}$ is chosen according to the distribution $P_{V|\mathbf{C},X_{i},U_{i},Z_{i}=l}$ except for $M^{\prime}_{iZ_{i}}=V$ which is Bob’s final view. Note that $P_{V|\mathbf{C},X_{1},X_{2},U_{1},U_{2},Z_{1},Z_{2}}=P_{V|\mathbf{C},X_{1},U_{1},Z_{1}}P_{V|\mathbf{C},X_{2},U_{2},Z_{2}}$ due to independence of transmitters. So, both $M^{\prime}_{i,l},l\in\{0,1\}$, have distribution $P_{V|\mathbf{C},X_{i},U_{i},Z_{i}=l}$. This makes the Markov chain $X^{\prime}_{i}\rightarrow(\mathbf{C},X_{i},U_{i})\rightarrow Z_{i}$, then we have $I(X^{\prime}_{i};Z_{i}|\mathbf{C},X_{i},U_{i})=0$, for $i\in\{1,2\}$.

### v4

@@ V-A2 OT over a point-to-point noisy channel
[Crepeau] A protocol $\Pi$ securely computes $\binom{2}{1}-\text{OT}^{k}$ perfectly if and only if for every pair of algorithms $\bar{A}=(\bar{A}_{1},\bar{A}_{2})$ that is admissible for protocol $\Pi$ and for all inputs $(X(m_{0},m_{1}),Z)$ and auxiliary input $\mathbf{C}$ (representing the total information transmitted over the noiseless public channel), $\bar{A}$ produces outputs $(U,V)$ such that the following conditions are satisfied:
@@
    $I(U^{\prime};Z^{\prime}|\mathbf{C},Z)=0,\,\,\,\text{and}\,\,\,I(U^{\prime};V|\mathbf{C},Z,Z^{\prime},M_{Z^{{}^{\prime}}})=0.$
@@
Consider a noisy DMC between Alice and Bob. OT with perfect secrecy (unconditional security) is impossible over a DM-MAC if one of the players is an unbounded cheater.
@@ V-A3 OT over a DM-MAC
    $I(X_{i};V|\mathbf{C},Z_{i},Z^{\prime}_{i},M_{iZ^{{}^{\prime}}_{i}})_{i\in\{1,2\}}$ $=0.$
@@ V-B OT over the DM-MAC
As previously discussed, from Alice’s perspective the SU–SBC is equivalent to supplying a uniform input to a binary MAC. In contrast, from Bob’s viewpoint the resulting observation can be represented as a distinguishable mixture of three channels: (i) a complete erasure channel with probability $(1-p)^{2}$, (ii) a partial erasure channel with probability 2p(1p)2p(1-p), and (iii) a DM–MAC with probability $p^{2}$.
@@
V-C Secret Key Agreement over MAC by Public Discussion
Secret key agreement over a simple point-to-point channel is initially addressed in [Maurer] and [Rudolf2]. In [Maurer], the author considers that Alice aims to share a secret key with Bob in the presence of a passive wiretapper. A noiseless one-way channel assists the main noisy channel with unlimited capacity. Also, it is assumed that Eve can receive all messages sent over the public channel without error, but the wiretapper can not change the messages without being detected.
A generalization of the above setting can be found in [Gohari1, Gohari2]. Also, the problem of secret key agreement over MAC is studied in [Aref], in which the authors assumed that each transmitter plays the role of wiretapper for the other transmitter. This is the MAC with confidential message [Poor]. Consider the following settings: Alice-1 and Alice-2 aim to share secret keys with Bob in the presence of a passive wiretapper. Alice-$i$, Bob, and Eve govern the input $X_{i}$, the output $Y$, and $E$, respectively. Furthermore, a public channel with unlimited capacity is available for one-way communication from the senders to Bob. Both resources (MAC and the public channel) are available for communication, but not at the same time. First, the senders are allowed to use the MAC $n$ times, and then they can use the public channel only once.
@@ VI-A Noiseless Binary Adder Channel (BAC)
Consider the following OT setting over the noiseless BAC [chou]: two senders, Alice-1 and Alice-2, each hold two binary strings. Bob must select one string from each sender, with the unchosen strings remaining hidden. Communication occurs via the BAC, where the channel output is given by the sum of the inputs, $Y=X_{1}+X_{2}$. This channel deterministically reveals the inputs when they are equal (0 or $1$), but produces an erasure symbol ($e$) when they differ, making the output ambiguous. Erasures thus occur in two of the four possible input pairs.
@@ VII Conclusion
We investigated bounds on the oblivious transfer (OT) capacity of noisy multiple-access channels (MACs), focusing on secure multiparty computation with two non-colluding senders and a single receiver. We proposed a protocol that is correct and secure against honest-but-curious parties and showed that it remains correct and secure even when the receiver is malicious, although the exact OT capacity in this adversarial setting remains open. For honest-but-curious players, we established the OT capacity region for a reduced MAC model, achieving optimal rates. For a malicious receiver, we characterized a feasible achievable rate region.
This work provides a comprehensive treatment of noisy MACs, introducing reductions to symmetric basic correlations (SBCs) as defined in [Winter1] and extending OT protocols to multi-sender configurations. Key contributions include a precise characterization of OT capacity in the honest-but-curious regime, the use of information-theoretic bounds to analyze performance, and significant progress toward understanding protocols in the presence of adversarial behavior.
An intriguing direction for future research is the study of OT over MACs with colluding parties. If one or both senders may collude with the receiver—a scenario more reflective of practical systems—the secrecy requirements become substantially more stringent. Moreover, in [Winter1], the authors present a protocol that remains correct and secure against a malicious sender under slightly unfair correlations, albeit without achieving positive rates. In the MAC setting, if both senders cheat, it is plausible that no positive OT rate can be achieved. However, the case in which only one sender is malicious remains largely unexplored. Investigating these challenging scenarios forms the foundation of our ongoing research efforts.
@@ Acknowledgment
The authors gratefully acknowledge valuable discussions with Holger Boche, Rémi A. Chou, Pin-Hsun Lin, and Andreas Winter during the course of this research.
CD and HA acknowledge financial support from the Federal Ministry of Research, Technology and Space (BMFTR) within the program Souverän. Digital. Vernetzt. as part of the joint project 6G-life2 (project ID 16KIS2415). They also gratefully acknowledge funding from the BMBF quantum programs QUIET (grant 16KISQ0170) and QD-CamNetz (grant 16KISQ169). Their research is further supported by the German Research Foundation (DFG) through the project Post-Shannon Theory and Implementation (grant DE 1915/2-1).
@@ Appendix B Proof of Lemma 2
Let $Z:=\sum_{i=0}^{n-1}Z_{i}$. Then $Z$ is a sum of $n$ i.i.d. bounded random variables in $0,\log(|\mathcal{X}|)]$. Hoeffding’s inequality [Hoeffding] states that for independent $Z_{i}\in[a,b]$:
@@ Appendix C Proof of Lemma 6
This lemma is an extended version of [Winter1, Lemma 5]. It evaluates the likelihood of passing a ”typicality test” when input sequences are fed into a noisy MAC and output sequences are generated. The lemma bounds the probability that the channel’s output remains consistent with a typical input-output relationship, even under deviations in the input strings (e.g., due to cheating). In simpler terms, this lemma demonstrates that if someone attempts to manipulate the inputs or outputs, the probability of such manipulations going undetected is exponentially small under specific conditions.
@@
Using the Chernoff bound [Winter1, Lemma 4] completes the proof. ∎
@@ Appendix D Proof of Theorem 9
$H(M_{\overline{Z^{{}^{\prime}}}}|V,Z,Z^{\prime},M_{Z^{{}^{\prime}}})=H(M_{\overline{Z^{{}^{\prime}}}}).$ (41)
@@
This implies that given the information $V$, the choice $Z$, the effective choice $Z^{\prime}$ and $M_{Z^{{}^{\prime}}}$, Bob has no additional information about $M_{\overline{Z^{{}^{\prime}}}}$. Consider the mutual information between Alice’s bits and the information received by Bob $I(X;V)\geq I(M_{0},M_{1};V)$. Since Bob should learn $M_{Z^{{}^{\prime}}}$ and nothing about $M_{\overline{Z^{{}^{\prime}}}}$, we have:
$I(M_{Z^{{}^{\prime}}},M_{\overline{Z^{{}^{\prime}}}};V|Z,Z^{\prime},M_{Z^{{}^{\prime}}})=$ $I(M_{Z^{{}^{\prime}}};V|Z,Z^{\prime},M_{Z^{{}^{\prime}}})+I(M_{\overline{Z^{{}^{\prime}}}};V|Z,Z^{\prime},M_{Z^{{}^{\prime}}}).$ (42)
@@
The first term should not be equal to zero because $V$ must also convey the value of $M_{Z^{{}^{\prime}}}$ to Bob. The second term should be zero due to the perfect secrecy of OT. So consider the second term:
$I(M_{\overline{Z^{{}^{\prime}}}};V|Z,Z^{\prime},M_{Z^{{}^{\prime}}})$ $=I(M_{\overline{Z^{{}^{\prime}}}};V|Z,Z^{\prime})$
$=H(M_{\overline{Z^{{}^{\prime}}}}|Z,Z^{\prime})-H(M_{\overline{Z^{{}^{\prime}}}}|V,Z,Z^{\prime})$
$=H(M_{\overline{Z^{{}^{\prime}}}}|Z,Z^{\prime})-H(M_{\overline{Z^{{}^{\prime}}}})$
$=-I(M_{\overline{Z^{{}^{\prime}}}};Z,Z^{\prime}),$ (43)
@@
where the first equality is happened due the fact that $M_{\overline{Z^{{}^{\prime}}}}$ and $M_{Z^{{}^{\prime}}}$ are independent, and the third equality follows from (41).
@@
As $I(M_{\overline{Z^{{}^{\prime}}}};Z,Z^{\prime})$ is a non-negative quantity, then we proved that $I(M_{\overline{Z^{{}^{\prime}}}};Y|M_{Z^{{}^{\prime}}},Z,Z^{\prime})$ is not a positive quantity. It is impossible unless $I(M_{\overline{Z^{{}^{\prime}}}};Z,Z^{\prime})=0$, which implies that all inputs $(X,Z)$ are independent (We already clarified that this is often unattainable and in the most general case, we assume there is a known dependency between the inputs). Combining (42) and (D), we have:
$I(X;V|Z,Z^{\prime},M_{Z^{{}^{\prime}}})$ $\stackrel{{\scriptstyle(a)}}{{\geq}}I(M_{Z^{{}^{\prime}}},M_{\overline{Z^{{}^{\prime}}}};V|Z,Z^{\prime},M_{Z^{{}^{\prime}}})$
$\stackrel{{\scriptstyle(b)}}{{=}}I(M_{Z^{{}^{\prime}}};V|Z,Z^{\prime},M_{Z^{{}^{\prime}}})$
$\,=H(M_{Z^{{}^{\prime}}}|Z,Z^{\prime},M_{Z^{{}^{\prime}}})-H(M_{Z^{{}^{\prime}}}|V,Z,Z^{\prime},M_{Z^{{}^{\prime}}})$
$\,=0,$ (44)
where $(a)$ is due to the Markov chain $(M_{0},M_{1})\to X\to V$ and data processing inequality (DPI), and $(b)$ is since given $V$ and $Z^{\prime}$, the uncertainty about $M_{Z^{{}^{\prime}}}$ is zero (correctness criterion). Also assume the usual OT correctness: given $V$ and the (effective) index $Z^{\prime}$, Bob recovers $M_{Z^{{}^{\prime}}}$ with certainty, so
$H(M_{Z^{{}^{\prime}}}\mid V,Z,Z^{\prime})=0.$ (45)
Finally assume the messages are independent a priori and non-degenerate:
$M_{0}\perp M_{1},\qquad H(M_{0})>0,\quad H(M_{1})>0.$
From the causal structure of the protocol we have the Markov chain $(M_{0},M_{1})\longrightarrow X\longrightarrow V$. Now consider (39):
$I(X;Z^{\prime}\mid Z)=0.$
Again use the Markov chain $(M_{0},M_{1})\to X\to Z^{\prime}$ (the variable $Z^{\prime}$ depends only on the transcript / view which in turn depends on $X$). By conditional DPI we get
$I(M_{0},M_{1};Z^{\prime}\mid Z)\leq I(X;Z^{\prime}\mid Z)=0,$
hence
$I(M_{0},M_{1};Z^{\prime}\mid Z)=0.$ (46)
Chain-rule on (46) gives
$0=I(M_{Z^{{}^{\prime}}};Z^{\prime}\mid Z)+I(M_{\overline{Z^{{}^{\prime}}}};Z^{\prime}\mid Z,M_{Z^{{}^{\prime}}}).$
Both terms are nonnegative, so each must be zero:
$I(M_{Z^{{}^{\prime}}};Z^{\prime}\mid Z)=0,\qquad I(M_{\overline{Z^{{}^{\prime}}}};Z^{\prime}\mid Z,M_{Z^{{}^{\prime}}})=0.$
From the first equality we deduce
$H(M_{Z^{{}^{\prime}}}\mid Z,Z^{\prime})=H(M_{Z^{{}^{\prime}}}\mid Z).$
(Conditional independence of $M_{Z}$ and $Z^{\prime}$ given $Z$ means conditioning on $Z^{\prime}$ does not further reduce entropy beyond conditioning on $Z$ alone.) By correctness (45) we have
$H(M_{Z^{{}^{\prime}}}\mid V,Z,Z^{\prime})=0.$
Combine this with (42)
$I(M_{0},M_{1};V\mid Z,Z^{\prime},M_{Z^{{}^{\prime}}})=0.$ (47)
The latter says that modulo $(Z,Z^{\prime},M_{Z^{{}^{\prime}}})$, $V$ is independent of $(M_{0},M_{1})$ — equivalently, all dependence of $V$ on the message space is already captured by the conditioned $M_{Z^{{}^{\prime}}}$. Concretely, we can rewrite (42) as
$H(V\mid Z,Z^{\prime},M_{Z^{{}^{\prime}}})=H(V\mid Z,Z^{\prime},M_{Z^{{}^{\prime}}},M_{0},M_{1}).$
But given $M_{0},M_{1}$ and the protocol, $V$ is determined stochastically through the protocol; the important observation is that (47) implies that, after conditioning on $(Z,Z^{\prime},M_{Z^{{}^{\prime}}})$, $V$ carries no additional information about the messages. Together with (45) this implies that the uncertainty of $M_{Z^{{}^{\prime}}}$ given $Z,Z^{\prime}$ must be no larger than its uncertainty after seeing $V$:
$H(M_{Z^{{}^{\prime}}}\mid Z,Z^{\prime})=H(M_{Z^{{}^{\prime}}}\mid V,Z,Z^{\prime})=0,$ (48)
where the last equality is due to (45).
This means that the message $M_{Z^{{}^{\prime}}}$ must be determined (with probability 1) by $(Z,Z^{\prime})$. Under the usual OT input model (two independent non-degenerate messages), this is impossible: $M_{Z^{{}^{\prime}}}$ has positive entropy conditioned on any small amount of side information unless the messages are degenerate (deterministic functions of the choices). Thus we get a contradiction to the assumption that $H(M_{0})>0$ or $H(M_{1})>0$. Therefore the perfect-information equalities (39) and (40) cannot hold simultaneously for nontrivial messages.∎
@@ Appendix E Proof of Theorem 10
The proof follows from the extended version presented in [Crepeau]. Consider the intuitive security criterion for Bob: if Alice-$i$ acts maliciously, she sends $X^{\prime}_{i}$ instead of $X_{i}$. This means that Alice-$i$ may replace any subset of bits by deterministic or random bits of her choice, so that the Hamming distance between the original and modified sequences satisfies $0\leq d_{H}\leq n$.
In such a case, the security constraint requires that the mutual information between Bob and Alice-$i$’s effective input, conditioned on $(X_{i},\mathbf{C})$, is zero $(X^{\prime}_{i}\rightarrow(\mathbf{C},X_{i})\rightarrow Y)$. Likewise, the mutual information between Bob’s final view $(V,Y)$ and Alice-$i$’s final view, conditioned on $(\mathbf{C},X_{i},X^{\prime}_{i})$, must also be zero $(U_{i}\rightarrow(\mathbf{C},X_{i},X^{\prime}_{i})\rightarrow(V,Y))$:
$I(Y;X^{\prime}_{i}\mid\mathbf{C},X_{i})=0,\qquad\text{and}\qquad I(V,Y;U_{i}\mid\mathbf{C},X_{i},X^{\prime}_{i})=0,$
where $Y=Z_{i}$ and $V=M^{\prime}_{iZ_{i}}$. Combining these conditions, we obtain
$I(Z_{i};X^{\prime}_{i}\mid$ $\mathbf{C},X_{i})+I(M^{\prime}_{iZ_{i}},Z_{i};U_{i}\mid\mathbf{C},X_{i},X^{\prime}_{i})$
$\,=I(Z_{i};X^{\prime}_{i}\mid\mathbf{C},X_{i})+I(Z_{i};U_{i}\mid\mathbf{C},X_{i},X^{\prime}_{i})+I(M^{\prime}_{iZ_{i}};U_{i}\mid\mathbf{C},Z_{i},X_{i},X^{\prime}_{i})$
$\stackrel{{\scriptstyle(a)}}{{=}}I(Z_{i};X^{\prime}_{i}\mid\mathbf{C},X_{i})+I(Z_{i};U_{i}\mid\mathbf{C},X_{i},X^{\prime}_{i})$
$\,=I(Z_{i};X^{\prime}_{i},U_{i}\mid\mathbf{C},X_{i})$
$\,=I(Z_{i};U_{i}\mid\mathbf{C},X_{i})+I(Z_{i};X^{\prime}_{i}\mid\mathbf{C},X_{i},U_{i})$
$\stackrel{{\scriptstyle(b)}}{{=}}I(Z_{i};U_{i}\mid\mathbf{C},X_{i})$
@@
where $(a)$ follows from $I(M^{\prime}_{iZ_{i}};U_{i}\mid\mathbf{C},Z_{i},X_{i},X^{\prime}_{i})=0$, and $(b)$ is justified as follows. Let $M^{\prime}_{iZ_{i}}=(M^{\prime}_{i,0},M^{\prime}_{i,1})$ where, for each $l\in\{0,1\}$, the value $M^{\prime}_{i,l}$ is chosen according to the distribution $P_{V\mid\mathbf{C},X_{i},U_{i},Z_{i}=l}$, except that $M^{\prime}_{iZ_{i}}=V$, which corresponds to Bob’s final view. Since
$P_{V\mid\mathbf{C},X_{1},X_{2},U_{1},U_{2},Z_{1},Z_{2}}=P_{V\mid\mathbf{C},X_{1},U_{1},Z_{1}}\,P_{V\mid\mathbf{C},X_{2},U_{2},Z_{2}}$
due to independence of the transmitters, both $M^{\prime}_{i,l}$ for $l\in\{0,1\}$ follow the distribution $P_{V\mid\mathbf{C},X_{i},U_{i},Z_{i}=l}$. This construction ensures the Markov chain
$X^{\prime}_{i}\rightarrow(\mathbf{C},X_{i},U_{i})\rightarrow Z_{i},$
which implies
$I(X^{\prime}_{i};Z_{i}\mid\mathbf{C},X_{i},U_{i})=0,$
for $i\in\{1,2\}$.